OpenStudy (pattipattt):

A train travels between two stations ½ mile apart in a minimum time of 41 sec. If the train accelerates and decelerates at 8 ft/sec^2, starting from rest at the first station and coming to a stop at the end of the station, what is its maximum speed in mph? how long does it travel at this top speed?

OpenStudy (phi):

V= a * t a is acceleration, 8 ft/sec^2 t is the time in seconds to reach max speed, which is half-way i.e. at t=41/2 seconds so \[ V= 8 \cdot \frac{41}{2} = 4\cdot 41 = 164 \text{ ft/second} \] Now you must change that to miles per hour